50.0 mL of 0.30 M potassium phosphate is mixed with 150.0 mL of 0.40 M magnesium phosphate. Determine the concentration of each ion in the final solution.
1
Expert's answer
2020-06-22T08:53:43-0400

Potassium phosphate is K3PO4 while Magnesium phosphate is Mg3(PO4)2

ionic reactions involved are

K3PO4 ----> 3K++PO43-

Mg3(PO4)2----> 3Mg2++ 2PO43-


no. of moles of K3PO4 = 50* 0.30= 15 mmoles

so according to stiochiometry in ionic reaction

no of moles of K+ = 3* 15 = 45 mmoles

while no. of PO43- = 15 mmoles


in case of magnesium phosphate

no of moles of Mg3(PO4)2 = 150* 0.40 = 60 mmoles

according to stiochiometry in ionic reaction

no of moles of Mg2+= 3* 60 = 180 mmoles

while for Phosphate ion = 2*60= 120 mmoles


total volume of solution = 50+ 150 mL = 150mL


concentration of K+= 45/200 = 0.225 M

concentration of Mg2+= 180/200= 0.90 M

concentration of PO43-=(15+120)/200

= 170/200 = 0.85 M